OpenStudy (anonymous):

(2^sin^2(3x))'= • A. ln8 × 2^sin^2(3x) × sin6x • B. ln2× 2^sin^2(3x) × cos3x • C. e^sin^2(3x) × ln2 × cos^2(3x) • D. 6cos3x × ln2 × 2 × sin^2(3x)

OpenStudy (anonymous):

Let y = 2^sin^2 (3x). Take the logarithm of both sides. log_2 y = log_2 2^sin^2 (3x) = sin^2 3x ln y / ln 2 = sin^2 3x ln y = ln 2 sin^2 3x Apply differentiation to both sides. y' / y = 6 ln 2 sin 3x cos 3x = 3 ln 2 sin 6x y' = 3y ln 2 sin 6x = 3 ln 2 sin 6x (2^sin^2 (3x)) Formulas and method used: Logarithmic differentiation log_a a^n = n log_b a = log_c a / log_c b 2 sin a cos a = sin (2a)

OpenStudy (anonymous):

Answer is A. 3 ln 2 sin 6x (2^sin^2 (3x)) = 2^sin^2 (3x) • ln 8 • sin 6x Another formula used a log_b c = log_b c^a

OpenStudy (anonymous):

tHANKS for the reply